y − 2x = −1
x + 3y = 4
one solution or no solution

Answer by flame8855(424)

y-2x = -1
y = 2x +1
subsitute
x+3(2x-1)= 4
x + 6x -3 = 4
7x = 7
x = 1
y = 2+1 = 3
(1,3)